OpenStudy (anonymous):

\[c \le ax+b \le d\] for x, if a < 0

OpenStudy (anonymous):

\[c \le ax+b \le d\] for x, if a < 0

OpenStudy (anonymous):

\[ c≤ax+b≤d \\ c -b \le a x \le d-b\\ \frac {c-b}a \ge x \ge \frac {d-b} a \]